OpenStudy (anonymous):

Perform a sequence of transformations to obtain the graph of each following functions. Do this step by step and explain each step along the way. 1. y=5/x 2.y=-4|-2x+3|

OpenStudy (anonymous):

Ok so you have to start with an elementary function and then work your way to the final function. Ill do the first for you let f(x)= 1/x scale f(x) by 5 f(x)*5 = 5/x y=5/x

OpenStudy (anonymous):

With 2. you will start with f(x) = abs(x)

OpenStudy (anonymous):

No, I meant like describe how to graph this... for example: 1. y = (x + 4)2 − 7 Step 1 we begin with y = x^2 and replace x with x + 4 to obtain y = (x + 4)^2. The graph of y=(x+ 4)^2 is a translation of the graph of y = x^2 to the left by 4 units. like help me explain the problems above.